# Beyond Scores: A Modular RAG-Based System for Automatic Short Answer Scoring with Feedback

### v1

@@ Abstract
Index Terms:
automatic short answer scoring, large language models, retrieval-augmented generation=-21pt
I Introduction
Feedback plays a critical role in student learning. Extensive research demonstrates that detailed and timely feedback significantly improves student performance by guiding learning and correcting misunderstandings [1, 2]. However, providing such elaborated feedback, especially for short answer questions, is a time-consuming and labor-intensive task for educators, particularly when managing large numbers of students. To address this, automatic short answer scoring (ASAS) systems have been developed to quickly assess student responses [3, 4]. Despite their utility, ASAS systems primarily focus on scoring correctness, offering holistic scores without the detailed feedback students need to understand their mistakes [5].
Current ASAS systems are typically trained on classification or regression tasks, where the objective is to assign a score to a student’s answer. While effective in evaluating answers, these systems lack the explainability needed for two key purposes: (1) helping students understand the reasoning behind their errors and the feedback given and (2) fostering trust in the scoring system [7, 6]. Without this granularity, students miss opportunities to improve, and educators may be reluctant to adopt automated grading systems.
In response to these limitations, recent research has focused on developing ASAS systems with feedback (ASAS-F) that not only score student answers but also generate detailed feedback [8, 9, 10]. These systems aim to provide students with actionable feedback that explains the reasoning behind their scores, helping them identify and correct their mistakes. Existing ASAS-F systems rely on resource-intensive fine-tuning with limited datasets, specific tasks [11, 12] or complex prompt engineering, all of which may not generalize well across contexts. Moreover, evaluating the quality of the textual feedback is usually only done using traditional statistical metrics, which do not capture the main aspects of quality, such as accuracy and clarity.
To overcome these challenges, we introduce a novel ASAS-F system that leverages large language models (LLMs) within a Retrieval-Augmented Generation (RAG) [13] framework. Inspired by similarity-based scoring [14], our method retrieves the most similar answers from a short answer scoring feedback dataset using the ColBERT retrieval [15] model. These retrieved answers serve as few-shot examples, allowing LLMs to generate both accurate scores and detailed feedback.
@@ I Introduction
Our experiments on the research questions demonstrate that the proposed approach significantly minimizes the need for extensive fine-tuning, resulting in a computationally efficient solution that maintains high accuracy while delivering clear, accurate feedback. The system is designed in a modular fashion, allowing it to adapt easily to various educational tasks without the need for extensive prompt engineering. By reducing reliance on large curated datasets, our method generates meaningful feedback with minimal adaptation, offering a scalable and cost-effective solution for educational applications. We share the code, model outputs and their evaluations from our experiments on Github to promote transparency and reproducibility 111Link available upon acceptance https://github.com/mennafateen/modular-asas-f-rag.
@@
II Related Work
II-A Automatic Short Answer Scoring
Automatic scoring has been studied extensively in the field of natural language processing (NLP) and educational technology [5, 16, 17, 18]. ASAS systems aim to grade students’ short answers automatically, offering immediate feedback, reducing teachers’ workload, and streamlining the grading process.
Traditional ASAS systems are often rule-based and rely on handcrafted features and heuristics to score short answers. While these systems are effective and have shown promising results on benchmark datasets, they require significant manual effort to develop [4, 19]. With the rise of deep learning, neural network-based and Transformer-based ASAS systems have been introduced, which automatically learn features from data and outperform traditional rule-based systems [3, 20].
Among the newer methods, instance-based approaches have gained traction. These approaches involve using pretrained models to map student responses directly to scores by pooling token representations across model layers [21]. These methods capture the context and semantics of input text, enabling them to outperform traditional rule-based systems. However, they often require large amounts of labeled data to achieve high performance and struggle in few-shot settings where data is limited. Comparisons between instance-based and similarity-based approaches have shown the latter to be more effective in zero-shot settings [14, 22, 23]. Additionally, recent advancements in dense information retrieval have proven effective in automatic essay scoring tasks [24]. Building on these insights, our work introduces a similarity-based approach using retrieval-augmented generation to enhance the performance of ASAS-F systems.
II-B ASAS Using Generative Models
@@
Few studies have explored the use of large language models (LLMs) for ASAS tasks. One study investigated the usage of LLMs for ASAS, focusing on zero-shot and few-shot settings across three diverse datasets [25]. The results showed that LLMs were able to achieve strong performance on tasks involving general knowledge questions but struggled with questions that required domain-specific knowledge or complicated reasoning.
Another study explored the use of proprietary LLMs, namely GPT3.5 and GPT4, for ASAS tasks [26], [27]. The findings indicated that while these models showed promise, their performance was negatively correlated with the length of student answers.
Additionally, research assessed the possibility of using ChatGPT (GPT-3.5) for automated grading [9]. The results showed a low correlation between the scores given by ChatGPT and human graders, with ChatGPT tending to give middle scores more frequently despite strong variation in student answers. Moreover, the study found that minor changes in answers could lead to significant differences in the scores given by ChatGPT.
II-C ASAS with Feedback
@@
The Short Answer Feedback (SAF) dataset [10] includes short answer questions in communication network topics, each with a reference answer, given score, label, and content-focused elaborated feedback. The SAF dataset provides a benchmark for evaluating ASAS systems that generate feedback, enabling the development of more informative and actionable systems. It also established a baseline for ASAS systems using the T5 model fine-tuned on the SAF dataset, demonstrating the feasibility of generating feedback for short answers.
A framework was proposed [8] that distils the rationale generation capability of ChatGPT by designing several prompt templates to generate feedback from ChatGPT, then fine-tuning a smaller language model on the refined generated outputs. The study showed that the fine-tuned model outperformed the original ChatGPT model on QWK scores.
Further research explored four different approaches for generating counterfactual feedback for short answers [28]. While some of the modified feedback could be graded as correct by automatic scoring systems, a domain expert deemed them incorrect.
Finally, one study explored the usage of LLMs for automated essay scoring (AES) with rationale [29], though the primary focus in ASAS is on content quality rather than writing style and structure. To the best of our knowledge, there is limited research evaluating LLMs for ASAS-F tasks, particularly in zero-shot and few-shot settings. Our work aims to bridge this gap by exploring the potential of LLMs for scoring short answers and generating feedback.
III Methodology
@@
III-A Problem Formulation
Traditional ASAS systems typically use classification or regression models to score student answers, focusing solely on assigning a numerical score. While effective for grading, these systems lack the ability to provide detailed feedback or explain the rationale behind a score, limiting their usefulness for student learning. In addition to such numeric scores and labels, ASAS-F systems additionally generate detailed elaborated feedback [1] that explains the reasoning behind the score.
@@ III-A Problem Formulation
III-B ASAS-F-Z: Zero-Shot ASAS-F
@@
To address these challenges, we utilize DSPy [30], a framework designed to automate prompt generation and refinement to realize our ASAS-F-Z and ASAS-F-RAG systems as shown in Figure 1.
@@
III-C ASAS-F-Opt: Automatic Few-Shot Optimization with DSPy
@@
We start by defining a performance metric for model evaluation. Then, we construct a Bayesian surrogate model that samples and assesses various combinations of prompts and examples. Our approach utilizes the MIPROv2 prompt optimization algorithm [31], ensuring that the most effective prompts and examples are identified to enhance scoring and feedback generation.
For simplicity, we use the accuracy of the generated labels as our optimization metric. We use the Llama3:70b [33] as the prompt generation model, and the smaller Mistral:7b [32] functions as the task executor. The prompt proposal model generates initial prompts based on the source code and analyzes the dataset to produce example traces. The optimizer then systematically selects and refines these examples based on their performance according to our defined metric. A code snippet of the implementation can be found in Appendix A.
III-D ASAS-F-RAG: Few-Shot ASAS-F Using RAG
@@
III-D1 Similarity-Based Majority-Vote with ColBERT
Many existing ASAS systems use similarity-based methods to assign scores to student answers [14]. These systems typically compare a student’s response to reference answers or high-scoring examples, relying on similarity metrics to determine the score.
@@ III-D1 Similarity-Based Majority-Vote with ColBERT
To implement this, we utilize the ColBERT model [15], which efficiently retrieves contextually relevant examples from the training data. Unlike traditional models that encode the entire input into a single vector, ColBERT encodes the input into a matrix of contextual token-level embeddings. This approach allows us to capture fine-grained similarities between the student’s answer and the training examples. Each word in the training example is represented by a BERT-based embedding, and these embeddings can be pre-stored for efficiency.
@@ III-D2 ASAS-F-RAG
IV Experimental Setup
IV-A Dataset
We evaluate our system on the Short Answer Feedback (SAF) Dataset [10]. The SAF dataset consists of short answer questions in communication network topics, each with a reference answer, given score, label and conent-focused elaborated feedback. There are no other datasets, based on our review, that provide such feedback for short answers. The dataset is split into training (70%), unseen answers (UA) (12%) and unseen questions (UA) (18%). The test split of UA includes new answers to the existing training questions, while the UQ split introduces novel questions. However, it is important to note that in our approach, both splits are considered novel or unseen questions, as we do not fine-tune or train the model on the training split. Even in the few-shot setting, we utilize less than 0.5% of the training data as the provided examples.
IV-B Evaluation Metrics
IV-B1 Scoring Metrics
@@ IV-B2 Feedback Evaluation
IV-C Model Selection
We employ various large language models, namely Mistral:7b, Mixtral:8x22b [32], Llama3:8b, and Llama3:70b [33] selected for their advanced performance and availability. We benchmark our ASAS-F system against several baselines: a majority class classifier, the T5 model [34] fine-tuned on the SAF dataset [10], and the Llama2:7b model fine-tuned on the SAF dataset [35]. The Llama2:7b finetune baseline is only available for the regression task, and has not been evaluated on the label classification task.
V Results
V-A Scoring Performance Analysis
V-A1 Similarity-Based Majority-Vote with ColBERT
@@
We evaluate the performance of the retrieval-based majority-vote classification approach on accuracy, F1 and RMSE. This serves as a preliminary experiment. Table I shows the performance of the ColBERT-based majority-vote classifier on the SAF dataset. The results indicate that the ColBERT-based majority-vote classifier outperforms the majority class baseline across all metrics showing the efficiency of using our refined approach.
V-A2 ASAS-F-Z
@@
Table II and Figures 4a and 4b show the results of the zero-shot ASAS-F system on both test splits of the SAF dataset compared to the baselines. In the UA split, the fine-tuned baselines outperform the zero-shot system in terms of accuracy and F1 score. However, in the UQ split, excluding the Llama3:8b model, the ASAS-F-Z system is able to outperform all baselines in all metrics. The highest performing model is the smallest Mistral:7b model, which achieves an accuracy of 0.755 and an F1 score of 0.751 in the UQ split. Llama3:70b achieves the lowest RMSE. This suggests that while finetuning can perform similarly to human raters in answers that it has seen during training, it tends to underperform when evaluating on new questions. The zero-shot ASAS-F system, on the other hand, is able to generalize better to new questions, indicating the potential of this approach for educational applications.
V-A3 ASAS-F-Opt
@@ V-A4 ASAS-F-RAG
To address the limitations shown by the ASAS-F-Z system especially in the UA split and the low performance shown by ASAS-F-Opt, we explore the ASAS-F-RAG system. This system focuses on automatically retrieving the most similar answers from existing data to augment the prompts used to improve model performance. Table III shows the results of the ASAS-F-RAG system on both test splits of the SAF dataset compared to the baselines. Figures 5a and 5b illustrate the performance of the ASAS-F-RAG system on the UA and UA splits, respectively.
@@
V-B Feedback Quality Analysis
V-B1 Statistical Analysis
Table IV presents the evaluation of feedback quality on the SAF dataset using SacreBLEU [36], ROUGE-2 [37], and BERTScore [38] metrics. It is evident that the baseline models were able to outperform both the zero-shot and few-shot ASAS-F systems in all metrics. However, the baseline models due to finetuning have been reported to often copy common phrases from the training data. Looking at the second-best performing models, we can see that the ASAS-F-RAG outperforms the ASAS-F-Z in both splits. This suggests that incorporating a small number of labeled examples can improve the quality of the generated feedback in terms of consistency with the reference feedback. However, increasing the number of examples beyond a certain threshold does not necessarily lead to higher similarity to the reference feedback. This can be seen in the UQ split where the second-best performing models used 3 examples.

### v2

@@ Abstract
Keywords automatic short answer scoring large language models retrieval-augmented generation
1 Introduction
Feedback plays a critical role in student learning. Extensive research demonstrates that detailed and timely feedback significantly improves student performance by guiding learning and correcting misunderstandings Shute (2008); Hattie and Timperley (2007). However, providing such elaborated feedback, especially for short answer questions, is a time-consuming and labor-intensive task for educators, particularly when managing large numbers of students. To address this, automatic short answer scoring (ASAS) systems have been developed to quickly assess student responses Riordan et al. (2017); Ramachandran et al. (2015). Despite their utility, ASAS systems primarily focus on scoring correctness, offering holistic scores without the detailed feedback students need to understand their mistakes Burrows et al. (2015).
Current ASAS systems are typically trained on classification or regression tasks, where the objective is to assign a score to a student’s answer. While effective in evaluating answers, these systems lack the explainability needed for two key purposes: (1) helping students understand the reasoning behind their errors and the feedback given and (2) fostering trust in the scoring system Winstone et al. (2017); Liu et al. (2018); Khosravi et al. (2022). Without this granularity, students miss opportunities to improve, and educators may be reluctant to adopt automated grading systems.
In response to these limitations, recent research has focused on developing ASAS systems with feedback (ASAS-F) that not only score student answers but also generate detailed feedback Li et al. (2023); Schneider et al. (2023); Filighera et al. (2022a). These systems aim to provide students with actionable feedback that explains the reasoning behind their scores, helping them identify and correct their mistakes. Existing ASAS-F systems rely on resource-intensive fine-tuning with limited datasets, specific tasks Dzikovska et al. (2014); Keuning et al. (2018) or complex prompt engineering, all of which may not generalize well across contexts. Moreover, evaluating the quality of the textual feedback is usually only done using traditional statistical metrics, which do not capture the main aspects of quality, such as accuracy and clarity.
To overcome these challenges, we introduce a novel ASAS-F system that leverages large language models (LLMs) within a Retrieval-Augmented Generation (RAG) Lewis et al. (2020) framework. Inspired by similarity-based scoring Bexte et al. (2023), our method retrieves the most similar answers from a short answer scoring feedback dataset using the ColBERT retrieval Khattab and Zaharia (2020) model. These retrieved answers serve as few-shot examples, allowing LLMs to generate both accurate scores and detailed feedback.
@@ 1 Introduction
Our experiments on the research questions demonstrate that the proposed approach significantly minimizes the need for extensive fine-tuning, resulting in a computationally efficient solution that maintains high accuracy while delivering clear, accurate feedback. The system is designed in a modular fashion, allowing it to adapt easily to various educational tasks without the need for extensive prompt engineering. By reducing reliance on large curated datasets, our method generates meaningful feedback with minimal adaptation, offering a scalable and cost-effective solution for educational applications. We share the code, model outputs and their evaluations from our experiments on Github to promote transparency and reproducibility 111Link available upon acceptance.
@@
2 Related Work
2.1 Automatic Short Answer Scoring
Automatic scoring has been studied extensively in the field of natural language processing (NLP) and educational technology Burrows et al. (2015); Xi (2010); Balfour (2013); Ke and Ng (2019). ASAS systems aim to grade students’ short answers automatically, offering immediate feedback, reducing teachers’ workload, and streamlining the grading process.
Traditional ASAS systems are often rule-based and rely on handcrafted features and heuristics to score short answers. While these systems are effective and have shown promising results on benchmark datasets, they require significant manual effort to develop Ramachandran et al. (2015); Kumar et al. (2019). With the rise of deep learning, neural network-based and Transformer-based ASAS systems have been introduced, which automatically learn features from data and outperform traditional rule-based systems Riordan et al. (2017); Sung et al. (2019).
Among the newer methods, instance-based approaches have gained traction. These approaches involve using pretrained models to map student responses directly to scores by pooling token representations across model layers Steimel and Riordan (2020). These methods capture the context and semantics of input text, enabling them to outperform traditional rule-based systems. However, they often require large amounts of labeled data to achieve high performance and struggle in few-shot settings where data is limited. Comparisons between instance-based and similarity-based approaches have shown the latter to be more effective in zero-shot settings Bexte et al. (2023); Wang et al. (2023); Fateen and Mine (2023). Additionally, recent advancements in dense information retrieval have proven effective in automatic essay scoring tasks Albatarni et al. (2024). Building on these insights, our work introduces a similarity-based approach using retrieval-augmented generation to enhance the performance of ASAS-F systems.
2.2 ASAS Using Generative Models
@@
Few studies have explored the use of large language models (LLMs) for ASAS tasks. One study investigated the usage of LLMs for ASAS, focusing on zero-shot and few-shot settings across three diverse datasets Chamieh et al. (2024). The results showed that LLMs were able to achieve strong performance on tasks involving general knowledge questions but struggled with questions that required domain-specific knowledge or complicated reasoning.
Another study explored the use of proprietary LLMs, namely GPT3.5 and GPT4, for ASAS tasks Chang and Ginter (2024), Achiam et al. (2023). The findings indicated that while these models showed promise, their performance was negatively correlated with the length of student answers.
In a recent study, researchers examined the use of ChatGPT for holistic essay scoring and compared its performance to human raters Tate et al. (2024). The findings showed that ChatGPT’s scoring was not statistically significantly different from human ratings and demonstrated substantial agreement. However, this study highlighted the need for further research particularly regarding improving scoring precision.
Additionally, research assessed the possibility of using ChatGPT (GPT-3.5) for automated grading Schneider et al. (2023). The results showed a low correlation between the scores given by ChatGPT and human graders, with ChatGPT tending to give middle scores more frequently despite strong variation in student answers. Moreover, the study found that minor changes in answers could lead to significant differences in the scores given by ChatGPT.
2.3 ASAS with Feedback
@@
The Short Answer Feedback (SAF) dataset Filighera et al. (2022a) includes short answer questions in communication network topics, each with a reference answer, given score, label, and content-focused elaborated feedback. The SAF dataset provides a benchmark for evaluating ASAS systems that generate feedback, enabling the development of more informative and actionable systems. It also established a baseline for ASAS systems using the T5 model fine-tuned on the SAF dataset, demonstrating the feasibility of generating feedback for short answers.
A framework was proposed Li et al. (2023) that distils the rationale generation capability of ChatGPT by designing several prompt templates to generate feedback from ChatGPT, then fine-tuning a smaller language model on the refined generated outputs. The study showed that the fine-tuned model outperformed the original ChatGPT model on QWK scores.
Further research explored four different approaches for generating counterfactual feedback for short answers Filighera et al. (2022b). While some of the modified feedback could be graded as correct by automatic scoring systems, a domain expert deemed them incorrect.
Finally, one study explored the usage of LLMs for automated essay scoring (AES) with rationale Yancey et al. (2023), though the primary focus in ASAS is on content quality rather than writing style and structure. To the best of our knowledge, there is limited research evaluating LLMs for ASAS-F tasks, particularly in zero-shot and few-shot settings. Our work aims to bridge this gap by exploring the potential of LLMs for scoring short answers and generating feedback.
3 Methodology
@@
3.1 Problem Formulation
Traditional ASAS systems typically use classification or regression models to score student answers, focusing solely on assigning a numerical score. While effective for grading, these systems lack the ability to provide detailed feedback or explain the rationale behind a score, limiting their usefulness for student learning. In addition to such numeric scores and labels, ASAS-F systems additionally generate detailed elaborated feedback Shute (2008) that explains the reasoning behind the score.
@@ 3.1 Problem Formulation
3.2 ASAS-F-Z: Zero-Shot ASAS-F
@@
To address these challenges, we utilize DSPy Khattab et al. (2023), a framework designed to automate prompt generation and refinement to realize our ASAS-F-Z and ASAS-F-RAG systems as shown in Figure 1.
@@
3.3 ASAS-F-Opt: Automatic Few-Shot Optimization with DSPy
@@
We start by defining a performance metric for model evaluation. Then, we construct a Bayesian surrogate model that samples and assesses various combinations of prompts and examples. Our approach utilizes the MIPROv2 prompt optimization algorithm Opsahl-Ong et al. (2024), ensuring that the most effective prompts and examples are identified to enhance scoring and feedback generation.
For simplicity, we use the accuracy of the generated labels as our optimization metric. We use the Llama3:70b Touvron et al. (2023) as the prompt generation model, and the smaller Mistral:7b Jiang et al. (2024) functions as the task executor. The prompt proposal model generates initial prompts based on the source code and analyzes the dataset to produce example traces. The optimizer then systematically selects and refines these examples based on their performance according to our defined metric. A code snippet of the implementation can be found in Appendix A.
3.4 ASAS-F-RAG: Few-Shot ASAS-F Using RAG
@@
3.4.1 Similarity-Based Majority-Vote with ColBERT
Many existing ASAS systems use similarity-based methods to assign scores to student answers Bexte et al. (2023). These systems typically compare a student’s response to reference answers or high-scoring examples, relying on similarity metrics to determine the score.
@@ 3.4.1 Similarity-Based Majority-Vote with ColBERT
To implement this, we utilize the ColBERT model Khattab and Zaharia (2020), which efficiently retrieves contextually relevant examples from the training data. Unlike traditional models that encode the entire input into a single vector, ColBERT encodes the input into a matrix of contextual token-level embeddings. This approach allows us to capture fine-grained similarities between the student’s answer and the training examples. Each word in the training example is represented by a BERT-based embedding, and these embeddings can be pre-stored for efficiency.
@@ 3.4.2 ASAS-F-RAG
4 Experimental Setup
4.1 Dataset
We evaluate our system on the Short Answer Feedback (SAF) Dataset Filighera et al. (2022a). The SAF dataset consists of short answer questions in communication network topics, each with a reference answer, given score, label and conent-focused elaborated feedback. There are no other datasets, based on our review, that provide such feedback for short answers. The dataset is split into training (70%), unseen answers (UA) (12%) and unseen questions (UA) (18%). The test split of UA includes new answers to the existing training questions, while the UQ split introduces novel questions. However, it is important to note that in our approach, both splits are considered novel or unseen questions, as we do not fine-tune or train the model on the training split. Even in the few-shot setting, we utilize less than 0.5% of the training data as the provided examples.
4.2 Evaluation Metrics
4.2.1 Scoring Metrics
@@ 4.2.2 Feedback Evaluation
4.3 Model Selection
We employ various large language models, namely Mistral:7b, Mixtral:8x22b Jiang et al. (2024), Llama3:8b, and Llama3:70b Touvron et al. (2023) selected for their advanced performance and availability. We benchmark our ASAS-F system against several baselines: a majority class classifier, the T5 model Raffel et al. (2020) fine-tuned on the SAF dataset Filighera et al. (2022a), and the Llama2:7b model fine-tuned on the SAF dataset Katuka et al. (2024). The Llama2:7b finetune baseline is only available for the regression task, and has not been evaluated on the label classification task.
5 Results
5.1 Scoring Performance Analysis
5.1.1 Similarity-Based Majority-Vote with ColBERT
@@
We evaluate the performance of the retrieval-based majority-vote classification approach on accuracy, F1 and RMSE. This serves as a preliminary experiment. Table 1 shows the performance of the ColBERT-based majority-vote classifier on the SAF dataset. The results indicate that the ColBERT-based majority-vote classifier outperforms the majority class baseline across all metrics showing the efficiency of using our refined approach.
5.1.2 ASAS-F-Z
@@
Table 2 and Figures 4a and 4b show the results of the zero-shot ASAS-F system on both test splits of the SAF dataset compared to the baselines. In the UA split, the fine-tuned baselines outperform the zero-shot system in terms of accuracy and F1 score. However, in the UQ split, excluding the Llama3:8b model, the ASAS-F-Z system is able to outperform all baselines in all metrics. The highest performing model is the smallest Mistral:7b model, which achieves an accuracy of 0.755 and an F1 score of 0.751 in the UQ split. Llama3:70b achieves the lowest RMSE. This suggests that while finetuning can perform similarly to human raters in answers that it has seen during training, it tends to underperform when evaluating on new questions. The zero-shot ASAS-F system, on the other hand, is able to generalize better to new questions, indicating the potential of this approach for educational applications.
5.1.3 ASAS-F-Opt
@@ 5.1.4 ASAS-F-RAG
To address the limitations shown by the ASAS-F-Z system especially in the UA split and the low performance shown by ASAS-F-Opt, we explore the ASAS-F-RAG system. This system focuses on automatically retrieving the most similar answers from existing data to augment the prompts used to improve model performance. Table 3 shows the results of the ASAS-F-RAG system on both test splits of the SAF dataset compared to the baselines. Figures 5a and 5b illustrate the performance of the ASAS-F-RAG system on the UA and UA splits, respectively.
@@
5.2 Feedback Quality Analysis
5.2.1 Statistical Analysis
Table 4 presents the evaluation of feedback quality on the SAF dataset using SacreBLEU Post (2018), ROUGE-2 Lin (2004), and BERTScore Zhang et al. (2019) metrics. It is evident that the baseline models were able to outperform both the zero-shot and few-shot ASAS-F systems in all metrics. However, the baseline models due to finetuning have been reported to often copy common phrases from the training data. Looking at the second-best performing models, we can see that the ASAS-F-RAG outperforms the ASAS-F-Z in both splits. This suggests that incorporating a small number of labeled examples can improve the quality of the generated feedback in terms of consistency with the reference feedback. However, increasing the number of examples beyond a certain threshold does not necessarily lead to higher similarity to the reference feedback. This can be seen in the UQ split where the second-best performing models used 3 examples.